OpenStudy (anonymous):

how to evaluate 8-(-2)?

OpenStudy (anonymous):

8-(-2) two negatives make a positive so 8+2=10

OpenStudy (anonymous):

8-(-2) = 8+(2) v 10

OpenStudy (anonymous):

how to evaluate (-3)-(-6)?

OpenStudy (anonymous):

-3-(-6)=-3+6=3

OpenStudy (anonymous):

8-(-2) = 8+(2) =10

OpenStudy (anonymous):

-3-(-6) =-3+6 =3

OpenStudy (anonymous):

how to evaluate (-5)+(-3)?

OpenStudy (anonymous):

-5-3=-8